Government spending in Robok is $150 billion, and its only tax is an income tax with a marginal tax rate of 0.3.

a. The balance on the government’s budget at a GDP level of $440 billion is a   (Click to select)   deficit   surplus  of $  billion.

b. The balance on the government’s budget at a GDP level of $550 billion is a   (Click to select)   deficit   surplus  of $  billion.

c. At what level of GDP will the economy of Robok have a balanced budget?


   Robok will have a balanced budget at a GDP level of $  billion.
Answer #1

a) government spending is 150 billion dollar. Government revenue from tax is 0.3"440 = 132 billion dollars. Therefore the balance on government budget is 132 - 150 = deficit of 18 billion dollars

b) government spending is 150 billion dollar. Government revenue from tax is 0.3"550 = 165 billion dollars. Therefore the balance on government budget is 165 - 150 = surplus of 15 billion dollars

c) required GDP level is

0.3*Y + 150 = 0

Y = 500 billion dollar

The answer is 500 billion dollar